# An asymptotic stability proof and a port-Hamiltonian physics-informed neural network approach to chaotic synchronization

### v2

@@ 3.1 Asymptotic stability proof of the synchronization manifold
$\Psi(e_{u};u_{1},u_{2},m,J)=\begin{cases}|2m-1|e_{u}^{2}+2|m|(J-1)|e_{u}|,&\text{if }u_{1}\geq 1,\;-1<u_{2}<1,\\ -e_{u}^{2}-4|m|e_{u},&\text{if }u_{1}\geq 1,\;u_{2}\leq-1,\\ -e_{u}^{2}-2|m|(J-1)e_{u},&\text{if }-1<u_{1}<1,\;u_{2}\geq 1,\\ |2m-1|e_{u}^{2},&\text{if }-1<u_{1}<1,\;-1<u_{2}<1,\\ -e_{u}^{2}-2|m|(J+1)e_{u},&\text{if }-1<u_{1}<1,\;u_{2}\leq-1,\\ -e_{u}^{2}+4|m|e_{u},&\text{if }u_{1}\leq-1,\;u_{2}\geq 1,\\ |2m-1|e_{u}^{2}+2|m|(J+1)e_{u},&\text{if }u_{1}\leq-1,\;-1<u_{2}<1,\\ -e_{u}^{2},&\text{otherwise.}\end{cases}$ (11)
The bound in Eq. (10) can be recast in matrix notation as
$\dot{V}\leq-\mathbf{E}^{\top}\mathbf{M}\mathbf{E}-|e_{u}|\Psi(e_{u};u_{1},u_{2},m,J),$ (12)
where $\mathbf{E}^{\top}=\left(|e_{x}|,|e_{y}|,|e_{z}|,|e_{u}|,|e_{\phi}|\right)$ denotes the vector of absolute error components and $\mathbf{M}$ is the symmetric matrix
$\mathbf{M}=\begin{bmatrix}A_{xx}&A_{xy}&A_{xz}&A_{xu}&A_{x\phi}\\ A_{yx}&1&0&0&0\\ A_{zx}&0&|r|&0&0\\ A_{ux}&0&0&A_{uu}&0\\ A_{\phi x}&0&0&0&|q|\end{bmatrix}$, (13)
with entries specified by
[…]
The fixed point $e^{*}_{i}=0$, $i=x,y,z,u,\phi$ of the error system Eq. (6) will be Lyapunov stable once the right-hand side of Eq. (12) is nonpositive. This is guaranteed whenever the quadratic form $\mathbf{E}^{\top}\mathbf{M}\mathbf{E}$ compensates all potentially positive contributions coming from the memristive term $\Psi(e_{u};u_{1},u_{2},m,J)$, which is ensured by sufficiently strong positive definiteness of $\mathbf{M}$.
Lemma 3.3 (Sylvester’s Criterion).
Let $A$ be a real symmetric matrix. Then $A$ is positive definite if and only if each of its leading principal minors is strictly positive.
Evaluating the leading principal minors of $\mathbf{M}$ yields
[…]
Choosing the parameters $\{a,b,d,f,h,k,m,q,r,s,\rho,g_{e}\}$ such that all minors in Eq. (15) are strictly positive ensures $\mathbf{M}\succ 0$, and hence $\dot{V}\leq 0$. In this situation, $V$ is positive definite and decrescent, so the fixed point $e^{*}_{i}=0$, $i=x,y,z,u,\phi$, of the error system Eq. (6) is Lyapunov stable.
@@
To strengthen this to global asymptotic stability of $e^{*}_{i}=0$, we must additionally show that errors decay to zero, i.e., that transverse perturbations to the synchronization manifold $\mathcal{M}$ vanish as $t\to\infty$. From Eqs. (7) and (12), the error components are uniformly bounded, $e_{i}\in L^{\infty}$ for $i=x,y,z,u,\phi$. Consequently, the right-hand side of Eq. (6) is also bounded, and thus $\dot{e_{i}}\in L^{\infty}$ for each component.
Next, decompose (12) as $\dot{V}\leq\dot{V}_{1}+\dot{V}_{2}$, where $\dot{V}_{1}=-\mathbf{e}^{\top}\mathbf{M}\mathbf{e}$ and $\dot{V}_{2}=-|e_{u}|\Psi(e_{u};u_{1},u_{2},m,J)$. Integrating the contribution from $\dot{V}_{1}$ yields
$\begin{split}\begin{aligned} \int_{0}^{t}\mathbf{e}(s)^{\top}\mathbf{M}\mathbf{e}(s)ds&=-\int_{0}^{t}\dot{V}_{1}(s)ds\\ &=V_{1}(0)-V_{1}(t)\leq V_{1}(0).\end{aligned}\end{split}$ (16)
@@
(Rayleigh bound). For any nonzero vector $x\in\mathbb{R}^{n}$ and any symmetric matrix $A\in\mathbb{R}^{n\times n}$, the quadratic form satisfies $x^{\top}Ax\geq\lambda_{\min}(A)\,\|x\|^{2},$ where $\lambda_{\min}(A)$ denotes the smallest eigenvalue of $A$.
Combining Eq. (16) with Lemma 3.4, we obtain the estimate
$\begin{split}\begin{aligned} \int_{0}^{t}\|\textbf{e}(s)\|ds\leq V_{1}(0)/\lambda_{min}(\mathbf{M}),\end{aligned}\end{split}$ (17)
where $\lambda_{min}(\mathbf{M})$ is the minimum eigenvalue of the positive definite matrix $\mathbf{M}$.
@@
From Eqs. (16)–(17) and Lemma 3.5, it follows that, for any initial condition,
$\begin{array}[]{lcl}\lim\limits_{t\rightarrow\infty}\|e_{i}(t)\|=0,i=x,y,z,u,\phi\\[5.69054pt] \end{array}.$ (19)
Therefore, the coupled chaotic systems in Eq. (2) synchronize globally and asymptotically in the sense of CS. ∎
Numerical simulations are provided in Sect. 4 to corroborate the effectiveness of the Theorem 3.2.
3.2 Synchronization energy
@@
According to Helmholtz’s theorem [54], any smooth vector field $\mathbf{F}(\mathbf{r})$ of a dynamical system can be decomposed into a conservative (divergence-free) part $\mathbf{F}_{c}(\mathbf{r})$ and a dissipative (curl-free) part $\mathbf{F}_{d}(\mathbf{r})$:
@@
where the conservative component governs rotational motion, while the dissipative component governs energy exchange with the environment. The divergence-free (conservative) and curl-free (dissipative) vector fields of the error dynamical system in Eq. (6) are respectively given by
@@
For the conservative vector field, Eq. (21) yields a first-order linear partial differential equation (PDE) for the Hamilton energy function:
@@
$H(e_{x},e_{y},e_{z},e_{u},e_{\phi})=2d\,x_{1}\,e_{x}^{2}+e_{y}^{2}-4dkf\,u_{1}x_{1}^{2}\,e_{u}^{2}-2d\rho\,x_{1}^{2}\,e_{\phi}^{2}+e_{z}-rs\,e_{\phi}+C,$ (24)
where $C$ is an arbitrary integration constant.
For the dissipative vector field, Eq. (21) gives the differential equation governing the time evolution of the Hamilton energy:
@@
$\quad-8dkf\,u_{1}x_{1}^{2}\,\beta(u_{1},u_{2})\,e_{u}+4d\rho q\,x_{1}^{2}\,e_{\phi}^{2}+qrs\,e_{\phi}.$
The synchronization energy function $H(e_{x},e_{y},e_{z},e_{u},e_{\phi})$ in Eq. (24) represents the instantaneous energy stored in the synchronization manifold, while its temporal derivative $dH/dt$ quantifies the energy flow associated with the synchronization process. A positive $dH/dt$ indicates net energy absorption (external energy input required to sustain synchrony), whereas a negative value corresponds to energy dissipation through coupling and intrinsic neuronal losses.
This formulation provides a quantitative measure of the energetic cost of synchronization: it defines the amount of energy per unit time required to maintain a given degree of synchrony. Consequently, variations in $H$ and $dH/dt$ under changes of system parameters reveal how coupling strength, dissipation, and intrinsic neuronal dynamics jointly determine both the stability and the energetic efficiency of the synchronized state [47]. The numerical simulations of $H$ and $dH/dt$ are provided in Sect. 4.

### v3

@@ 3.1 Asymptotic and practical stability in the linearized error dynamical system
$\widehat{\Psi}(e_{u};u_{1},u_{2},m,J)\mathrel{\mathop{\ordinarycolon}}=\begin{cases}\Big(|2m-1|+\dfrac{\varepsilon}{2}\Big)e_{u}^{2}+\dfrac{2|m|^{2}(J-1)^{2}}{\varepsilon},&\text{if }u_{1}\geq 1,\;-1<u_{2}<1,\\[4.0pt] -e_{u}^{2},&\text{if }u_{1}\geq 1,\;u_{2}\leq-1,\\[4.0pt] -e_{u}^{2},&\text{if }-1<u_{1}<1,\;u_{2}\geq 1,\\[4.0pt] |2m-1|\,e_{u}^{2},&\text{if }-1<u_{1}<1,\;-1<u_{2}<1,\\[4.0pt] -e_{u}^{2},&\text{if }-1<u_{1}<1,\;u_{2}\leq-1,\\[4.0pt] -\Big(1-\dfrac{\varepsilon}{2}\Big)e_{u}^{2}+\dfrac{8|m|^{2}}{\varepsilon},&\text{if }u_{1}\leq-1,\;u_{2}\geq 1,\\[4.0pt] \Big(|2m-1|+\dfrac{\varepsilon}{2}\Big)e_{u}^{2}+\dfrac{2|m|^{2}(J+1)^{2}}{\varepsilon},&\text{if }u_{1}\leq-1,\;-1<u_{2}<1,\\[4.0pt] -e_{u}^{2},&\text{otherwise.}\end{cases}$ (14)
Under the strict dissipativity assumption of the theorem, we define
$\widehat{\widehat{\Psi}}(e_{u};u_{1},u_{2},m,J)\mathrel{\mathop{\ordinarycolon}}=\widehat{\Psi}(e_{u};u_{1},u_{2},m,J)+e_{u}^{2},$ (15)
so that $\widehat{\widehat{\Psi}}(e_{u};u_{1}(t),u_{2}(t),m,J)\leq 0$ for all $t\geq 0$ in the dissipative Cases $2,3,5,8$ in Eq. (14). Substituting all these bounds into Eq. (10) and collecting terms yields the compact semi-quadratic-form estimate
$\dot{V}\leq-\mathbf{e}^{\top}\mathbf{M}\mathbf{e}+\widehat{\widehat{\Psi}}(e_{u};u_{1},u_{2},m,J).$ (16)
In particular, compared to the quadratic-form bound obtained directly from Eq. (10), the additional $e_{u}^{2}$ dissipation is absorbed into the quadratic form, which increases the $e_{u}^{2}$-weight by $1$. And $\mathbf{M}$ here is the diagonal and hence symmetric matrix
$\mathbf{M}=\begin{bmatrix}A_{xx}&0&0&0&0\\ 0&A_{yy}&0&0&0\\ 0&0&A_{zz}&0&0\\ 0&0&0&A_{uu}&0\\ 0&0&0&0&A_{\phi\phi}\end{bmatrix}$, (17)
with the nonzero (diagonal) entries given by
$\left\{\begin{aligned} A_{xx}&=-(3|a|+2|kf|)J^{2}-\left(2|b|+|d|+\frac{3}{2}|\rho|\right)J-|kh|+2|g_{e}|-\frac{1}{2}|rs|-\frac{3}{2},\\ A_{yy}&=\frac{1}{2}-|d|J,\\ A_{zz}&=|r|-\frac{1}{2}|rs|,\\ A_{uu}&=\frac{1}{2}-|kf|J^{2},\\ A_{\phi\phi}&=|q|-\frac{1}{2}|\rho|J-\frac{1}{2}.\end{aligned}\right$. (18)
Since $\mathbf{M}$ is diagonal, $\mathbf{M}\succ 0$ if and only if
$A_{xx}>0,\qquad A_{yy}>0,\qquad A_{zz}>0,\qquad A_{uu}>0,\qquad A_{\phi\phi}>0,$ (19)
which can be achieved with suitable parameter values of $\{a,b,d,f,h,k,m,q,r,s,\rho,g_{e},J\}$. The fixed point $\mathbf{e}\equiv 0$ of the linearized error system in Eq. (6) is Lyapunov stable once the right-hand side of Eq. (16) is nonpositive, i.e., $\dot{V}\leq 0$. To ensure this, we impose the dissipativity assumption that the trajectory remains in memristive regimes for which
$\widehat{\Psi}(e_{u};u_{1}(t),u_{2}(t),m,J)\leq-e_{u}^{2},\qquad\forall t\geq 0,$ (20)
for all values of $e_{u}$ (i.e., Cases $2,3,5,8$ in Eq. (14)). Thus, under Eqs. (19) and (20), the inequality in Eq. (16) reduces to
$\dot{V}\leq-\,\mathbf{e}^{\mathsf{T}}\mathbf{M}\mathbf{e}+\widehat{\widehat{\Psi}}(e_{u};u_{1}(t),u_{2}(t),m,J)\leq-\,\mathbf{e}^{\mathsf{T}}\mathbf{M}\mathbf{e}\leq 0,$ (21)
so that $V(t)$ is nonincreasing; since $V(e)=\tfrac{1}{2}\|\mathbf{e}\|^{2}$ is positive definite, this implies that $\mathbf{e}\equiv 0$ is Lyapunov stable.
@@
To strengthen this Lyapunov stability to asymptotic stability, we must additionally show that the errors decay to zero as $t\to\infty$, i.e., that transverse perturbations to the synchronization manifold vanish as $t\to\infty$. From Eq. (7) and the bounds in Eqs. (16) and (14), the error components are uniformly bounded, i.e., $e_{i}\in L^{\infty}$ for $i=x,y,z,u,\phi$. Since $x_{1}(t),u_{1}(t),\phi_{1}(t)$ are bounded, all time-varying coefficients appearing in Eq. (6) are bounded functions of time. Moreover, the additive terms in the piecewise definition of $\dot{e}_{u}$ (e.g., $\pm 4m$ and $2m(u_{1}\pm 1)$) are bounded because $u_{1}$ is bounded. Hence the right-hand side of Eq. (6) is bounded whenever $\mathbf{e}(t)$ is bounded, and therefore $\dot{\mathbf{e}}(t)\in L^{\infty}$ (equivalently, each component $\dot{e}_{i}\in L^{\infty}$).
@@
(Rayleigh bound). For any nonzero vector $x\in\mathbb{R}^{n}$ and any symmetric matrix $A\in\mathbb{R}^{n\times n}$, the quadratic form satisfies $x^{\top}Ax\geq\lambda_{\min}(A)\,\|x\|^{2},$ where $\lambda_{\min}(A)$ is the smallest eigenvalue of $A$.
With a positive definite symmetric $\mathbf{M}$, Lemma 3.4 gives
$\bm{e}^{\top}\mathbf{M}\bm{e}\geq\lambda_{\min}(\mathbf{M})\,\|\bm{e}\|^{2},$ (22)
where $\lambda_{\min}(\mathbf{M})>0$ is the minimum eigenvalue of $\mathbf{M}$. From Eq. (21) and the Rayleigh bound Eq. (22), we obtain for all $t\geq 0$,
$\dot{V}(t)\leq-\,\mathbf{e}(t)^{\mathsf{T}}\mathbf{M}\mathbf{e}(t)\leq-\lambda_{\min}(\mathbf{M})\,\|\mathbf{e}(t)\|^{2},$ (23)
where $\lambda_{\min}(\mathbf{M})>0$ denotes the smallest eigenvalue of $\mathbf{M}$. Integrating Eq. (23) from 0 to $t$ yields
$V(t)-V(0)\leq-\lambda_{\min}(\mathbf{M})\int_{0}^{t}\|\mathbf{e}(s)\|^{2}\,ds,$ (24)
and since $V(t)\geq 0$ for all $t\geq 0$, it follows that
$\lambda_{\min}(\mathbf{M})\int_{0}^{t}\|\mathbf{e}(s)\|^{2}\,ds\leq V(0)-V(t)\leq V(0).$ (25)
Hence we obtain the uniform bound
$\int_{0}^{t}\|\mathbf{e}(s)\|^{2}\,ds\leq\frac{V(0)}{\lambda_{\min}(\mathbf{M})}<\infty,\qquad\forall\,t\geq 0,$ (26)
and therefore $\|\mathbf{e}(\cdot)\|^{2}\in L^{1}([0,\infty))$.
@@
To apply Barbalat’s lemma, we define $f(t)=\|\mathbf{e}(t)\|^{2}$. From Eq. (26), $f\in L^{1}([0,\infty))$, i.e., $\int_{0}^{\infty}f(s)\,ds<\infty$. Since $e_{i}\in L^{\infty}$ and $\dot{e}_{i}\in L^{\infty}$ for each component $(i=x,y,z,u,\phi)$, we have that
$\dot{f}(t)=\frac{d}{dt}\|\mathbf{e}(t)\|^{2}=2\mathbf{e}(t)^{\top}\dot{\mathbf{e}}(t),$ (27)
is also bounded. Hence $f(t)=\|\mathbf{e}(t)\|^{2}$ is Lipschitz and thus uniformly continuous on $0,\infty)$. By Barbalat’s lemma, $f(t)\to 0$ as $t\to\infty$, i.e.,
$\lim\limits_{t\rightarrow\infty}\|\mathbf{e}(t)\|=0,$ (28)
which establishes asymptotic transverse stability of the synchronization manifold for the linearized error dynamics of Eq. (6).
In Theorem 3.3 we obtain asymptotic convergence $\|\mathbf{e}(t)\|\to 0$, which establishes complete synchronization in the sense of Definition 3.1.
Theorem 3.6 (Practical stability under non-dissipative memristive contribution).
Consider the linearized transverse error dynamics in Eq. (6) along a bounded reference trajectory and assume that there exists $J>0$ such that $|x_{1}(t)|,|u_{1}(t)|,|\phi_{1}(t)|\leq J$ for all $t\geq 0$. Let $V(\mathbf{e})=\tfrac{1}{2}\|\mathbf{e}\|^{2}$ and let $\mathbf{M}$ be the (diagonal) matrix in Eq. (18), and assume that $\mathbf{M}\succ 0$. Fix any $\varepsilon\in(0,2)$ and let $\widehat{\Psi}$ be the Young upper bound in Eq. (14).
Assume that, for all $t\geq 0$, the memristive switching remains in the non-dissipative Cases $1,4,6,7$ of Eq. (14). Define
$\gamma\mathrel{\mathop{\ordinarycolon}}=|2m-1|+\frac{\varepsilon}{2},\qquad C\mathrel{\mathop{\ordinarycolon}}=\frac{1}{\varepsilon}\max\Big\{2|m|^{2}(J-1)^{2},\;8|m|^{2},\;2|m|^{2}(J+1)^{2}\Big\}.$ (29)
If $\alpha\mathrel{\mathop{\ordinarycolon}}=\lambda_{\min}(\mathbf{M})-\gamma>0$ (with $\lambda_{\min}(\mathbf{M})=\min\{A_{xx},A_{yy},A_{zz},A_{uu},A_{\phi\phi}\}$), then along solutions of Eq. (6) one has
$\dot{V}(t)\leq-2\alpha V(t)+C,\qquad t\geq 0.$ (30)
Consequently,
$V(t)\leq\Big(V(0)-\frac{C}{2\alpha}\Big)e^{-2\alpha t}+\frac{C}{2\alpha},\qquad t\geq 0,$ (31)
and in particular
$\limsup_{t\to\infty}\|\mathbf{e}(t)\|\leq\sqrt{\frac{C}{\alpha}}.$ (32)
Thus the origin is practically stable and solutions ultimately enter and remain in the ball $\|\mathbf{e}\|\leq\sqrt{C/\alpha}$.
Proof.
From Eq. (16) we have
$\dot{V}(t)\leq-\,\mathbf{e}(t)^{\mathsf{T}}\mathbf{M}\,\mathbf{e}(t)+\widehat{\Psi}\big(e_{u}(t);u_{1}(t),u_{2}(t),m,J\big).$ (33)
Since $\mathbf{M}\succ 0$ is symmetric, the Rayleigh bound yields
$\mathbf{e}^{\mathsf{T}}\mathbf{M}\,\mathbf{e}\geq\lambda_{\min}(\mathbf{M})\,\|\mathbf{e}\|^{2}.$ (34)
Next, assume the switching remains in non-dissipative Cases $1,4,6,7$ of Eq. (14). We claim that for all $t\geq 0$,
$\widehat{\Psi}\big(e_{u}(t);\cdot\big)\leq\gamma\,e_{u}(t)^{2}+C,$ (35)
with $\gamma$ and $C$ given in Eq. (29). Indeed:
  • In Case $4$, $\widehat{\Psi}=|2m-1|e_{u}^{2}\leq\gamma e_{u}^{2}$.
  • In Case $1$, $\widehat{\Psi}=\big(|2m-1|+\frac{\varepsilon}{2}\big)e_{u}^{2}+\frac{2|m|^{2}(J-1)^{2}}{\varepsilon}\leq\gamma e_{u}^{2}+C$.
  • In Case $6$, $\widehat{\Psi}=-(1-\frac{\varepsilon}{2})e_{u}^{2}+\frac{8|m|^{2}}{\varepsilon}\leq\frac{8|m|^{2}}{\varepsilon}\leq C\leq\gamma e_{u}^{2}+C$, since $\varepsilon\in(0,2)$.
  • In Case $7$, $\widehat{\Psi}=\big(|2m-1|+\frac{\varepsilon}{2}\big)e_{u}^{2}+\frac{2|m|^{2}(J+1)^{2}}{\varepsilon}\leq\gamma e_{u}^{2}+C$.
This proves Eq. (35). Since $C$ is chosen as the maximum of the constants appearing in Cases $1,6,7$, the bound holds uniformly in time whenever the switching remains in Cases $1,4,6,7$. Using $e_{u}^{2}\leq\|\mathbf{e}\|^{2}$, we obtain
$\widehat{\Psi}\big(e_{u}(t);\cdot\big)\leq\gamma\|\mathbf{e}(t)\|^{2}+C.$
Substituting this and Eq. (34) into Eq. (33) yields
$\dot{V}(t)\leq-\lambda_{\min}(\mathbf{M})\|\mathbf{e}(t)\|^{2}+\gamma\|\mathbf{e}(t)\|^{2}+C=-\alpha\|\mathbf{e}(t)\|^{2}+C,$
where $\alpha=\lambda_{\min}(\mathbf{M})-\gamma>0$ by assumption. Since $\|\mathbf{e}\|^{2}=2V$, we obtain Eq. (30):
$\dot{V}(t)\leq-2\alpha V(t)+C.$
Let $W$ solve $\dot{W}=-2\alpha W+C$ with $W(0)=V(0)$. Then
$W(t)=\Big(V(0)-\frac{C}{2\alpha}\Big)e^{-2\alpha t}+\frac{C}{2\alpha}.$
By the Comparison Principle, $V(t)\leq W(t)$ for all $t\geq 0$, which proves Eq. (31). Taking $t\to\infty$ gives $\limsup_{t\to\infty}V(t)\leq\frac{C}{2\alpha}$, hence
$\limsup_{t\to\infty}\|\mathbf{e}(t)\|^{2}=2\limsup_{t\to\infty}V(t)\leq\frac{C}{\alpha},$
which yields Eq. (32) and completes the proof.
In Theorem 3.6 we obtain an explicit ultimate bound, which establishes practical ($\delta$-) synchronization in the sense of Definition 3.2, yielding $\delta=\sqrt{C/\alpha}$, which can be arbitrary small.
Remark 2 (Asymptotic versus practical stability under memristive switching).
The distinction between Theorem 3.3 and Theorem 3.6 is driven by the sign structure of the memristive contribution. In the dissipative regimes (Cases $2,3,5,8$), one has $\widehat{\Psi}\leq-e_{u}^{2}$ along the trajectory, so the Lyapunov estimate strengthens to $\dot{V}\leq-\mathbf{e}^{\mathsf{T}}\mathbf{M}\mathbf{e}-e_{u}^{2}\leq-\mathbf{e}^{\mathsf{T}}\mathbf{M}\mathbf{e}$, which implies $\|\mathbf{e}(t)\|\to 0$ as $t\to\infty$. In contrast, in the remaining switching regimes (Cases $1,4,6,7$) the Young upper bound $\widehat{\Psi}$ can contain a nonnegative quadratic term in $e_{u}^{2}$ and/or a positive constant offset. In this case the best general estimate is of the form $\dot{V}\leq-2\alpha V+C$ with $C>0$, yielding uniform boundedness and ultimate convergence of the error to a ball whose radius scales like $\sqrt{C/\alpha}$, rather than asymptotic convergence to the origin unless the offset vanishes (i.e., $C=0$, which can occur only when the memristive-feedback state-switching nonlinearity parameter $m\in[0,1]$ is exactly zero).
3.2 Synchronization Hamiltonian and its rate identity
@@
According to Helmholtz’s theorem [55], any smooth vector field $\mathbf{F}(\mathbf{r})$ of a dynamical system can be decomposed into a conservative (divergence-free) part $\mathbf{F}_{c}(\mathbf{r})$ and a dissipative (curl-free) part $\mathbf{F}_{d}(\mathbf{r})$ as follow:
@@
where the conservative component governs rotational motion, while the dissipative component governs the synchronization Hamiltonian exchange with the environment. The divergence-free (conservative) and curl-free (dissipative) vector fields of the linearized error dynamical system in Eq. (6) are respectively given by
@@
For the conservative vector field, Eq. (37) yields a first-order linear partial differential equation for the Hamiltonian function:
@@
$H(e_{x},e_{y},e_{z},e_{u},e_{\phi})=2d\,x_{1}\,e_{x}^{2}+e_{y}^{2}-4dkf\,u_{1}x_{1}^{2}\,e_{u}^{2}-2d\rho\,x_{1}^{2}\,e_{\phi}^{2}+e_{z}-rs\,e_{\phi}+K,$ (40)
where $K$ is an arbitrary integration constant.
Remark 3.
$H$ is defined up to an additive constant $K$ (gauge freedom, which we fix by setting $H(0)=0$), and is not guaranteed to be positive definite or bounded below (as it includes linear terms (e.g., $e_{z}$, $rse_{\phi}$, and an arbitrary constant $K$) as a function on $\mathbb{R}^{5}$. Thus, $H$ is scalar potential associated with the conservative part of the error field, and $\dot{H}=dH/dt$ gives a rate identity.
For the dissipative vector field, Eq. (37) gives the differential equation governing the time evolution of the synchronization Hamiltonian :
@@
$-8dkf\,u_{1}x_{1}^{2}\,\beta(u_{1},u_{2})\,e_{u}+4d\rho q\,x_{1}^{2}\,e_{\phi}^{2}+qrs\,e_{\phi}.$
$H(e_{x},e_{y},e_{z},e_{u},e_{\phi})$ represents a synchronization Hamiltonian potential (defined up to a constant), while it temporal derivative $dH/dt$ quantifies the instantaneous rate of change of this potential along the error dynamics and can be used as an energetic diagnostic of synchronization transients
This formulation provides a quantitative measure of the energetic cost of synchronization: it defines the amount of energy per unit time required to maintain a given degree of synchrony. Consequently, variations in $H$ and $dH/dt$ under changes of system parameters reveal how coupling strength, dissipation, and intrinsic neuronal dynamics jointly determine both the stability and the energetic efficiency of the synchronized state [48]. The numerical simulations of $H$ and $dH/dt$ are provided in Section 4.